A new elementary school is to have 50 faculty members. Some will be teachers and the rest will be aides. At this school, teachers a will earn an average of $20,000 a year and aides will earn an average of $12,000 a year. If the total amount spent is $896,000 on salaries, how many teachers and aides will there be?
Answer by jorel1380(3719)

Let t be the amount of teachers. Then the amount of aides would be 50-t. So:
20000t+12000(50-t)=896000
20t+12(50-t)=896
8t=896-600=296
t=37
There will be 37 teachers and 13 aides. ☺☺☺